write in polar form: 2sqroot2(1-i)
Found 2 solutions by fractalier, venugopalramana:
Answer by fractalier(6550)

r = sqrt(x^2 + y^2)
theta = arctan(y/x) ...I use q for theta
so that
r = sqrt(8 + 8) = 4 and
q = arctan (-1) = 315° or -45°
so we have in polar form
(4, 315°)

Answer by venugopalramana(3286)

write in polar form: 2sqroot2(1-i)
SEE THE FOLLOWING AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK
write in polar form: negative sqroot of 5+i
Z=5+I
|Z|=R=SQRT(5^2+1^2)=SQRT(26)
TAN(THETA)=1/5.....THETA=TAN^-1 (1/5)=A SAY
Z=R*E^(iA)={(26)^0.5}*{E^(iA)}
-SQRT(Z)=-{(26)^0.25}*E^(iA/2)=-2.258*E^(0.1974i)....ANGLE IN RADIANS ..OR...
=-2.258*E^(11.316i)....ANGLE IN DEGREES